# Efficient Large Language Models: A Survey

### v2

@@ 2.3.2. Memory-Efficient Fine-Tuning
Speculative Decoding (Leviathan
et al., 2022),
@@
Staged Speculative (Spector and Re, 2023),
BiLD (Kim et al., 2023d),
SpecInfer (Miao et al., 2023),
LLMA (Yang et al., 2023a),
Medusa (Cai
et al., 2023),
Santilli et al. (2023), leaf, text width=23em
@@
SkipDecode (Corro et al., 2023),
Heavy Hitter Oracle ()
(Zhang et al., 2023d),
Dynamic Context Pruning (Anagnostidis et al., 2023), Scissorhands (Liu et al., 2023a), leaf, text width=22em
@@
Flash-Decoding (Dao
et al., 2023),
FlashDecoding++ (Hong et al., 2023),leaf, text width=31em
2.4. Efficient Inference
Efficient inference aims to enhance the efficiency of the inference process for LLMs. As summarized in Figure 10, efficient inference techniques can be grouped into techniques at the algorithm level and system level.
@@ 2.4. Efficient Inference
Speculative Decoding. Speculative decoding (i.e., speculative sampling) (Leviathan et al., 2022) is a decoding strategy for autoregressive language models that speed up sampling by parallel token computation through using smaller draft models to create speculative prefixes for the larger target model. Chen et al. (2023a) propose to run a faster autoregressive model times and then evaluate the preliminary output with the large target LLM. A tailored rejection sampling strategy is employed to approve a selection of the draft tokens in a left-to-right order, thereby recapturing the distribution of the target model during the procedure. Staged Speculative (Spector and Re, 2023) transforms the speculative batch into a tree structure representing potential token sequences. This restructuring aims to expedite the generation of larger and improved speculative batches. It introduces an additional phase for speculative decoding of the initial model, thereby enhancing overall performance. BiLD (Kim et al., 2023d) optimizes speculative decoding through two innovative strategies: the fallback policy that permits the smaller draft model to waive control to the larger target model when it lacks sufficient confidence, and the rollback policy that enables the target model to revisit and rectify any inaccurate predictions made by the smaller draft model. SpecInfer (Miao et al., 2023) speeds up inference by employing speculative inference techniques and token tree validation. Its core idea involves merging a range of small speculative models that have been fine-tuned collectively to collaboratively forecast the output of the LLM, which is then used to validate all the predictions. LLMA (Yang et al., 2023a) chooses a text segment from a closely related reference and duplicates its tokens into the decoder. It then concurrently assesses the suitability of these tokens as the decoding output within a single decoding step. This approach results in a speed increase of more than two times for LLMs while maintaining the same generated results as traditional greedy decoding. Medusa (Cai et al., 2023) involves freezing the LLM backbone, fine-tuning additional heads, and using a tree-based attention mechanism to process predictions in parallel to speed up the decoding process. Lastly, Santilli et al. (2023) propose parallel decoding including the Jacobi and Gauss-Seidel fixed-point iteration methods for speculative decoding. Among these strategies, Jacobi decoding was extended into Lookahead decoding (Fu et al., 2023a) to enhance the efficiency of LLMs.
KV-Cache Optimization. Minimizing the repeated computation of Key-Value (KV) pairs during the inference process of LLMs is also key to enhancing the inference efficiency. Corro et al. (2023) propose SkipDecode, a token-level early exit approach that utilizes a unique exit point for each token in a batch at every sequence position, and skips the lower and middle layers to accelerate the inference process. Zhang et al. (2023d) point out that KV-cache is scaling linearly with the sequence length and batch size. They propose a KV cache eviction strategy that formulates the KV cache eviction as a dynamic sub-modular problem and dynamically retains a balance between recent and important tokens, reducing the latency for LLMs inference. Dynamic Context Pruning (Anagnostidis et al., 2023) utilizes a learnable mechanism to identify and remove non-informative KV-cache tokens. In doing so, it not only enhances efficiency but also improves interpretability. Liu et al. (2023a) underscore the Persistence of Importance Hypothesis, suggesting that only tokens that were crucial at an earlier phase will have a significant impact on subsequent stages. Based on this theory, they propose Scissorhands that introduces a streamlined algorithm for LLM inference using a compact KV-cache.
@@
System-Level Inference Efficiency Optimization. The efficiency of LLM inference can also be optimized at the system level. For example, FlexGen (Sheng et al., 2023) is a high-throughput inference engine that enables the execution of LLMs on GPUs with limited memory. It uses a linear programming-based search approach to coordinate various hardware, combining the memory and computation from GPU, CPU, and disk. Furthermore, FlexGen quantizes the weights and attention cache to 4 bits, increasing the inference speed of OPT-175B (Zhang et al., 2022d) on a single 16GB GPU. Deja Vu (Liu et al., 2023f) presents the notion of contextual sparsity, which is a collection of MLP and attention modules that produce the same result as a dense model, but with fewer components. This technique trains predictors to identify the sparsity and then uses kernel fusion and memory coalescing to speed up the inference process. Pope et al. (2023) develop a simple analytical framework to select the best multi-dimensional partitioning methods optimized for TPU v4 slices based on the application requirements. By combining this with some existing low-level optimizations, they have achieved greater efficiency on PaLM (Chowdhery et al., 2022) in comparison to the FasterTransformer (NVIDIA, 2023) standards. S (Jin et al., 2023) has created a system that is aware of the output sequence beforehand. It can anticipate the length of the sequence and arrange generation requests accordingly, optimizing the utilization of device resources and increasing the rate of production. Orca (Yu et al., 2022) employs iteration-level scheduling to decide batch sizes. When a sequence in a batch is completed, it is substituted with a new one, resulting in improved GPU utilization compared to static batching. DeepSpeed-Inference (Aminabadi et al., 2022) is a multi-GPU inference approach that is designed to enhance the efficiency of both dense and sparse Transformer models when they are contained within the collective GPU memory. Furthermore, it provides a mixed inference technique that utilizes CPU and NVMe memory, in addition to GPU memory and computation, guaranteeing high-throughput inference even for models that are too large to fit in the combined GPU memory. Flash-Decoding (Dao et al., 2023) is a technique that boosts the speed of long-context inference by breaking down keys/values into smaller pieces, computing attention on these pieces in parallel, and then combining them to generate the final output. FlashDecoding++ (Hong et al., 2023) supports mainstream language models and hardware backends through asynchronous softmax, double buffering for flat GEMM optimization, and heuristic dataflow, resulting in up to 4.86x and 2.18x acceleration on NVIDIA and AMD GPUs respectively compared to HuggingFace implementations.
@@
GQA (Ainslie et al., 2023), leaf, text width=9em
Feature Information Reduction, fill=blue!10
Funnel-Transformer (Dai
et al., 2020),
Nyströmformer (Xiong et al., 2021),
Set Transformer (Lee
et al., 2019), leaf, text width=30em
@@
Sumformer (Alberti
et al., 2023),
FluRKA (Gupta
et al., 2023),
@@
Linformer (Wang
et al., 2020), leaf, text width=33em
@@
Longformer (Beltagy
et al., 2020), Blockwise Transformer (Qiu
et al., 2019), Sparse Transformer (Child
et al., 2019), leaf, text width=31em
@@
HyperAttention (Han et al., 2023),
Reformer (Kitaev
et al., 2020),
Sparse Sinkhorn Attention (Tay
et al., 2020),
Clustered Attention (Vyas
et al., 2020),
ClusterFormer (Wang et al., 2022a),
Routing Transformer (Roy
et al., 2021), leaf, text width=32em
@@
FlashAttention (Dao
et al., 2022),
FlashAttention-2 (Dao, 2023),
PagedAttention (Kwon et al., 2023),
A (Ham et al., 2020),
ELSA (Ham
et al., 2021), leaf, text width=36em
@@
Switch Transformer (Fedus
et al., 2022),
@@
PanGu- (Ren
et al., 2023b),
Mixtral 8x7B (Jiang et al., 2023b), leaf, text width=25em
@@
Lifelong-MoE (Chen et al., 2023j),
Flan-MoE (Shen et al., 2023), leaf, text width=25em
System-Level MoE Optimization, fill=blue!10
FastMoE (He
et al., 2021),
FasterMoE (He
et al., 2022),
@@
SmartMoE (Zhai
et al., 2023),
MegaBlocks (Gale
et al., 2023), leaf, text width=37em
@@
Extrapolation and Interpolation, fill=blue!10
ALiBi (Press
et al., 2022),
xPOS (Sun et al., 2022a),
CLEX (Chen
et al., 2023d),
Linear Interpolation (Chen
et al., 2023h),
@@
YaRN Interpolation (Peng
et al., 2023c),
FIRE (Li
et al., 2023e),
PoSE (Zhu
et al., 2023)
, leaf, text width=36em
@@
Memformer (Wu
et al., 2020),
-former (Martins
et al., 2022),
RMT (Bulatov
et al., 2022),
@@
Retentive Network (Sun et al., [n.d.])
, leaf, text width=29em
@@
Mistral (Jiang et al., 2023b),
StreamingLLM (Xiao
et al., 2023b),
@@
SLED (Ivgi
et al., 2023)
, leaf, text width=31em
Memory-Retrieval Augmentation, fill=blue!10
KNN-augmented Transformer (Wu
et al., 2022),
Landmark Attention (Mohtashami and
Jaggi, 2023),
LongMem (Wang et al., 2023a),
Unlimiformer (Bertsch
et al., 2023),
@@
Xu et al. (2023a), leaf, text width=34em
Transformer-Alternative Architecture, fill=blue!10
@@
Diagonal State Space (Gupta
et al., 2022),
H3 (Fu
et al., 2023b),
Gated State Space (Mehta
et al., 2022),
@@
Mamba (Gu and Dao, 2023),
SMA (Ren
et al., 2023a), leaf, text width=37em
@@
MEGABYTE (Yu et al., 2023), leaf, text width=20em
2.5. Efficient Architecture Design
@@
Efficient architecture design for LLMs refers to the strategic optimization of model architecture and computational processes to enhance performance and scalability while minimizing resource consumption. Figure 12 summarizes efficient architecture designs for LLMs.
2.5.1. Efficient Attention
@@
The quadratic time and space complexity of attention modules considerably slows down the pre-training, inference and fine-tuning of LLMs (Keles et al., 2022). A lot of techniques have been proposed to make attention lightweight for more efficient execution. These techniques can be generally categorized as sharing-based attention, feature information reduction, kernelization or low-rank, fixed pattern strategies, learnable pattern strategies, and hardware-assisted attention.
@@
Sharing-based Attention. Sharing-based attention aims to accelerate attention computation during inference through different KV heads sharing schemes. For example, LLaMA-2 (Touvron et al., 2023b) optimizes the autoregressive decoding process by using multi-query attention (MQA) (Shazeer, 2019) and grouped-query attention (GQA) (Ainslie et al., 2023). In contrast to multi-head attention, which uses several attention layers (heads) simultaneously with distinct linear transformations for queries, keys, values, and outputs, MQA has all its heads sharing one set of keys and values. While MQA utilizes only one key-value head to speed up decoder inference, it might compromise quality. To address this, GQA offers a modified version of MQA by employing more than one key-value heads but fewer than the total number of query heads to enhance the inference quality.
Feature Information Reduction. The principle of feature information reduction, as evidenced by models such as Funnel-Transformer (Dai et al., 2020), Nyströmformer (Xiong et al., 2021), and Set Transformer (Lee et al., 2019), is to cut computation demands by reducing feature information within a sequence, which leads to a proportionate reduction in required computation resources. For example, Funnel-Transformer (Dai et al., 2020) reduces the sequence length of hidden states to decrease computational costs, while its decoder can reconstruct deep representations for each token from this compressed sequence.
@@
Kernelization or Low-Rank. Kernelization or low-rank techniques adopted by models such as Sumformer (Alberti et al., 2023), FluRKA (Gupta et al., 2023), Scatterbrain (Chen et al., 2021a), Low-Rank Transformer (LRT) (Winata et al., 2020), Performer (Choromanski et al., 2021), Random Feature Attention (RFA) (Peng et al., 2021), Linear Transformer (Katharopoulos et al., 2020), and Linformer (Wang et al., 2020), , enhance computational efficacy by utilizing low-rank representations of the self-attention matrix or by adopting attention kernelization techniques. Specifically, low-rank methods focus on compacting the dimensions of attention keys and values. For example, Linformer (Wang et al., 2020) proposes to segment scaled dot-product attention into smaller units via linear projection. Kernelization, a variant of low-rank technique, focuses on approximating the attention matrix (Choromanski et al., 2020). For example, Performer (Choromanski et al., 2021) condenses softmax attention-kernels using positive orthogonal random features. Sumformer (Alberti et al., 2023) approximates the equivariant sequence-to-sequence function, offering a universal solution for both Linformer and Performer.
@@
Fixed Pattern Strategies. Fixed pattern strategies adopted by models such as (Pagliardini et al., 2023), Big Bird (Zaheer et al., 2020), Poolingformer (Zhang et al., 2021), Longformer (Beltagy et al., 2020), Blockwise Transformer (Qiu et al., 2019), and Sparse Transformer (Child et al., 2019) improve efficiency by sparsifying the attention matrix. This is achieved by confining the attention scope to predetermined patterns, such as local windows or fixed-stride block patterns. For instance, Longformer (Beltagy et al., 2020)’s attention mechanism, designed as an alternative to conventional self-attention, merges local windowed attention with globally oriented attention tailored to specific tasks. Pagliardini et al. (2023) have expanded FlashAttention (Dao et al., 2022) to support a broad spectrum of attention sparsity patterns, including key-query dropping and hashing-based attention techniques.
@@
Learnable Pattern Strategies. Learnable pattern strategies adopted by models such as HyperAttention (Han et al., 2023), Reformer (Kitaev et al., 2020), Sparse Sinkhorn Attention (Tay et al., 2020), Clustered Attention (Vyas et al., 2020), ClusterFormer (Wang et al., 2022a), and Routing Transformer (Roy et al., 2021) improve efficiency by learning token relevance and subsequently grouping tokens into buckets or clusters. As an example, HyperAttention (Han et al., 2023) proposes a parameterization for spectral approximation and employs two key metrics: the maximal column norm in the normalized attention matrix and the row norm ratio in the unnormalized matrix after large entry removal. It also utilizes the learnable sort locality-sensitive hashing (sortLSH) technique and fast matrix multiplication via row norm sampling. Their experiment results show that HyperAttention enhances both inference and training speeds for LLMs with only minimal performance degradation.
Hardware-Assisted Attention. Besides algorithmic approaches that sparsify attentions and thereby streamline the computation of the attention matrix, several studies concentrate on realizing efficient and lightweight attention mechanisms from hardware aspects. For example, FlashAttention (Dao et al., 2022) and FlashAttention-2 (Dao, 2023) aim to reduce the communication times between GPU high-bandwidth memory (HBM) and GPU on-chip SRAM when calculating the attention module in LLMs. Instead of transmitting the values and results between HBM and SRAM multiple times as is done in the standard attention mechanism, FlashAttention combines all the attention operations into one kernel and tiles the weight matrices into smaller blocks to better fit the small SRAM. As a result, only one communication is required to process each attention block, significantly increasing the efficiency for processing the entire attention block. Inspired by virtual memory and paging techniques, PagedAttention (Kwon et al., 2023) enables the storage of continuous keys and values in non-contiguous memory space. Specifically, PagedAttention divides the KV cache of each sequence into blocks, each containing the keys and values for a fixed number of tokens. During the attention computation, the PagedAttention kernel manages these blocks efficiently by maintaining a block table to reduce memory fragmentation. Specifically, the contiguous logical blocks of a sequence are mapped to non-contiguous physical blocks via the table and the table automatically allocates a new physical block for every newly generated token. This reduces the amount of memory wasted when generating new tokens, thus improving its efficiency. A (Ham et al., 2020) introduces an innovative candidate selection process that reduces the number of keys and offers a custom hardware pipeline that taps into parallelism to speed up approximated attention techniques, further enhancing their efficiency. ELSA (Ham et al., 2021) uses Kronecker decomposition to approximate the attention module, which not only reduces its complexity, but also makes it more suitable for parallelization on hardware, thus making it more efficient when used for inference.
2.5.2. Mixture of Experts (MoE)

### v4

@@ 2.4 Efficient Inference
Speculative Decoding (Leviathan et al., 2023),
@@
Staged Speculative (Spector & Re, 2023),
BiLD (Kim et al., 2023c),
SpecInfer (Miao et al., 2024),
LLMA (Yang et al., 2023b),
Medusa (Cai et al., 2024),
Santilli et al. (2023), PaSS (Monea et al., 2023), leaf, text width=33em
@@
KIVI (Zirui Liu et al., 2023), KVQuant (Hooper et al., 2024),
Heavy-Hitter Oracle (H2O)
(Zhang et al., 2023f),
Scissorhands (Liu et al., 2023e), StreamingLLM (Xiao et al., 2024), leaf, text width=30em
@@
Flash-Decoding (Dao et al., 2023),
FlashDecoding++ (Hong et al., 2023),leaf, text width=53.5em
@@
Speculative Decoding. Speculative decoding (i.e., speculative sampling) (Leviathan et al., 2023) is a decoding strategy for autoregressive language models that speeds up the sampling process by computing tokens using a smaller draft model in parallel to create speculative prefixes for the large target model. Chen et al. (2023a) focus on the distributed serving setting for LLMs and propose to run a faster autoregressive model times and then evaluate the preliminary output with the large target model. A tailored rejection sampling strategy is employed to approve a selection of the draft tokens in a left-to-right order, thereby recapturing the distribution of the large target model during the procedure. Staged Speculative (Spector & Re, 2023) transforms the speculative batch into a tree structure representing potential token sequences. This restructuring expedites the generation of larger and improved speculative batches. It also introduces an additional phase for speculative decoding of the initial model, thereby enhancing overall performance, showing 1.36x over standard speculative decoding. BiLD (Kim et al., 2023c) optimizes speculative decoding through two innovative techniques: the fallback policy that permits the smaller draft model to waive control to the larger target model when it lacks sufficient confidence; and the rollback policy that enables the target model to revisit and rectify any inaccurate predictions made by the smaller draft model. SpecInfer (Miao et al., 2024) extends speculative decoding and speeds up inference by employing speculative inference techniques and token tree validation. Its core idea involves merging a range of small speculative models that have been fine-tuned collectively to collaboratively forecast the output of the large target model, which is then used to validate all the predictions. Different from speculative decoding that needs to introduce an additional efficient drafter model to generate a draft for checking, LLMA (Yang et al., 2023b) chooses a text segment from a closely related reference and duplicates its tokens into the decoder. It then concurrently assesses the suitability of these tokens as the decoding output within a single decoding step. This approach results in a speed increase of more than two times while maintaining the same generated results as traditional greedy decoding. Similarly, instead of using a separate draft model to sequentially generate candidate output, Medusa (Cai et al., 2024) proposes to freeze the LLM backbone, fine-tune additional heads, and use a tree-based attention mechanism to process predictions in parallel to speed up the decoding process. Lastly, Santilli et al. (2023) propose parallel decoding including the Jacobi and Gauss-Seidel fixed-point iteration methods for speculative decoding. Among these methods, Jacobi decoding was extended into Lookahead decoding (Fu et al., 2023c) to further enhance the efficiency.
KV-Cache Optimization. During inference, LLMs need to store the Key-Value (KV) pairs of the past tokens into the cache for future token generation. The size of KV cache needed enlarges massively with the increase of generated token length, resulting in considerable memory consumption and long inference latency. Therefore, reducing the size of KV cache is key to enhancing inference efficiency. Existing KV-cache optimization techniques can in general be grouped into two categories. The first category is to compress the KV cache. For example, Zirui Liu et al. (2023) propose KIVI, a tuning-free 2bit KV cache quantization algorithm which quantizes the key cache per-channel and the value cache per-token, achieving 2.6x less peak memory usage during inference. Similarly, Hooper et al. (2024) conduct an empirical study on the impact of per-channel quantization and other types of quantization such as quantization before rotary positional embedding. Based on their findings, they propose KVQuant which combines these quantization methods to quantize the KV cache of LLaMA to 3-bit. The second category of KV-Cache optimization techniques is to evict some KVs from the cache. For instance, Zhang et al. (2023f) propose Heavy-Hitter Oracle (H2O), a KV cache eviction strategy that formulates the KV cache eviction as a dynamic submodular problem and dynamically retains a balance between recent and performance-critical tokens, improving the throughput for LLMs inference. Similarly, Liu et al. (2023e) propose a hypothesis named the persistence of importance, suggesting that only tokens that were crucial at an earlier phase will have a significant impact on subsequent stages. Based on this hypothesis, they design Scissorhands which significantly reduces the KV cache without compromising model quality. Lastly, StreamingLLM (Xiao et al., 2024) incorporates window attention, where only the most recent KVs are cached into a fixed-size sliding window, into their algorithm design. Through evicting the outdated KVs, StreamingLLM ensures constant memory usage and decoding speed after the cache is initially filled.
@@
System-Level Inference Efficiency Optimization. The efficiency of LLM inference can also be optimized at the system level under a specific hardware architecture. For example, FlexGen (Sheng et al., 2023) is a high-throughput inference engine that enables the execution of LLMs on GPUs with limited memory. It uses a linear programming-based search approach to coordinate various hardware, combining the memory and computation from GPU, CPU, and disk. Furthermore, FlexGen quantizes the weights and attention cache to 4 bits, which increases the inference speed of OPT-175B (Zhang et al., 2022a) on a single 16GB GPU. Pope et al. (2023) develop a simple analytical framework to partition a model in order to scale Transformer inference based on the application requirements. By combining it with scheduling and memory optimizations, they are able to achieve better efficiency on PaLM (Chowdhery et al., 2022) in comparison to FasterTransformer (NVIDIA, 2023a). Orca (Yu et al., 2022) employs iteration-level scheduling to serve batched sequences with variable output sequence length. When a sequence in a batch is completed, it is returned to the user so that a new sequence can be served immediately. As a result, Orca improves GPU utilization compared to static batching, showing 36.9× throughput improvement under the same level of latency compared to FasterTransformer. S3 (Jin et al., 2023) creates a system that is aware of the output sequence beforehand. It can anticipate the length of the sequence and arrange generation requests accordingly, optimizing the utilization of device resources and increasing the rate of production, showing higher throughput than Orca with the same number of GPUs. However, both Orca and S3 lead to memory fragmentation due to their inaccurate memory provisioning for each request. vLLM (Kwon et al., 2023) addresses the memory efficiency problem with PagedAttention, which enables the storage of continuous keys and values in non-contiguous memory space. Specifically, PagedAttention divides the KV cache of each sequence into blocks, each containing the keys and values for a fixed number of tokens. During attention computation, PagedAttention kernel manages these blocks efficiently by maintaining a block table to reduce memory fragmentation. Specifically, the contiguous logical blocks of a sequence are mapped to non-contiguous physical blocks via the table and the table automatically allocates a new physical block for every newly generated token. This reduces the amount of memory wasted when generating new tokens, thus improving its efficiency, showing that PagedAttention improves the throughput of popular LLMs by 2-4× with the same level of latency compared to FasterTransformer (NVIDIA, 2023a) and Orca (Yu et al., 2022). On the other hand, infinitely optimizing server-side aggregated metrics does not necessarily lead to good user experience or Quality of Experience (QoE) especially under high server load. Andes (Liu et al., 2024b) first defines QoE for the LLM-based text streaming services and proposes a QoE-aware serving systems to optimize QoE by prioritizing requests based on their resource demand and service acquired. DeepSpeed-Inference (Aminabadi et al., 2022) is a multi-GPU inference approach designed to enhance the efficiency of both dense and sparse Transformer models when they are contained within the collective GPU memory. Furthermore, it provides a mixed inference technique that utilizes CPU and NVMe memory, in addition to GPU memory and computation, enabling high-throughput inference even for models that are too large to fit in the combined GPU memory. This approach demonstrates lower latency than FasterTransformer under the same throughput. Flash-Decoding (Dao et al., 2023) boosts the speed of long-context inference by breaking down keys/values into smaller pieces, computing attention on these pieces in parallel, and then combining them to generate the final output. It outperforms FasterTransformer and FlashAttention (Dao et al., 2022) in decoding speed for very large sequences. FlashDecoding++ (Hong et al., 2023) supports mainstream language models and hardware backends through asynchronous softmax, double buffering for flat GEMM optimization, and heuristic dataflow, resulting in up to 4.86x and 2.18x acceleration on Nvidia and AMD GPUs respectively compared to HuggingFace implementations, showing higher speedup compared to Flash-Decoding under the same throughput.
@@
GQA (Ainslie et al., 2023), leaf, text width=22em
@@
Sumformer (Alberti et al., 2023),
FluRKA (Gupta et al., 2023),
@@
Linformer (Wang et al., 2020), leaf, text width=43em
@@
Longformer (Beltagy et al., 2020), Blockwise Transformer (Qiu et al., 2020), Sparse Transformer (Child et al., 2019),
Lightning Attention-2 (Qin et al., 2024), leaf, text width=52em
@@
HyperAttention (Han et al., 2024),
Reformer (Kitaev et al., 2020),
Sparse Sinkhorn Attention (Tay et al., 2020),
Clustered Attention (Vyas et al., 2020),
ClusterFormer (Wang et al., 2022b),
Routing Transformer (Roy et al., 2021), leaf, text width=52em
@@
FlashAttention (Dao et al., 2022),
vAttention (Prabhu et al., 2024), leaf, text width=30em
@@
Switch Transformer (Fedus et al., 2022),
@@
PanGu- (Ren et al., 2023b),
Mixtral 8x7B (Jiang et al., 2023a), leaf, text width=44em
@@
Lifelong-MoE (Chen et al., 2023f),
Flan-MoE (Shen et al., 2024), leaf, text width=41em
System-Level MoE Optimization, fill=blue!10, text width=17em
FastMoE (He et al., 2021),
FasterMoE (He et al., 2022a),
@@
SmartMoE (Zhai et al., 2023),
MegaBlocks (Gale et al., 2023), leaf, text width=45em
@@
Positional Extrapolation and Interpolation, fill=blue!10, text width=24em
ALiBi (Press et al., 2022),
xPOS (Sun et al., 2023c),
CLEX (Chen et al., 2024a),
RoPE-PI (Chen et al., 2023d),
@@
YaRN Interpolation (Peng et al., 2024),
FIRE (Li et al., 2024a),
PoSE (Zhu et al., 2024), leaf, text width=40em
@@
Memformer (Wu et al., 2022a),
-former (Martins et al., 2022),
RMT (Bulatov et al., 2022),
@@
Retentive Network (Sun et al., 2023b)
, leaf, text width=54em
@@
Mistral (Jiang et al., 2023a),
StreamingLLM (Xiao et al., 2024),
@@
SLED (Ivgi et al., 2023), MemWalker (Chen et al., 2023c),
RAPTOR (Sarthi et al., 2024)
, leaf, text width=41em
Memory-Retrieval Augmentation, fill=blue!10,text width=21em
Memorizing Transformer (Wu et al., 2022c),
Landmark Attention (Mohtashami & Jaggi, 2023),
LongMem (Wang et al., 2023e),
Unlimiformer (Bertsch et al., 2023),
@@
Xu et al. (2024a), leaf, text width=42em
Transformer-Alternative Architecture, text width=20em, fill=blue!10
@@
Diagonal State Space (Gupta et al., 2022),
H3 (Fu et al., 2023a),
Gated State Space (Mehta et al., 2023),
@@
Mamba (Gu & Dao, 2023),
SMA (Ren et al., 2023a), leaf, text width=48em
@@
MEGABYTE (YU et al., 2023), leaf, text width=38em
2.5 Efficient Architecture Design
@@
Efficient architecture design for LLMs refers to the strategic optimization of model architecture and computational processes to enhance performance and scalability while minimizing resource consumption. Figure 12 provides a summary of existing efforts on designing efficient architectures for LLMs.
2.5.1 Efficient Attention
@@
The quadratic time and space complexity of attention modules considerably slows down the pre-training, inference, and fine-tuning of LLMs (Duman Keles et al., 2023). Many techniques have been proposed to make attention lightweight for more efficient execution. These techniques can be generally categorized as sharing-based attention, kernelization or low-rank, fixed pattern strategies, learnable pattern strategies, and hardware-assisted attention.
@@
Sharing-based Attention. Sharing-based attention accelerates attention computation during inference through KV heads sharing. For example, LLaMA-2 (Touvron et al., 2023b) optimizes the autoregressive decoding process by using multi-query attention (MQA) (Shazeer, 2019) and grouped-query attention (GQA) (Ainslie et al., 2023). In traditional multi-head attention (MHA), each head has distinct linear transformations for input matrix queries (Q), keys (K) and values (V), allowing for diverse representations and attention mechanisms across different subspaces. In MQA, all heads share a single set of key and value weights across all query heads. Thus, MQA speeds up the inference but could compromise output quality. To address this drawback, GQA interpolates MQA and MHA by employing one key and value heads for each group of query heads to enhance inference quality.
@@
Kernelization or Low-Rank. Kernelization or low-rank techniques adopted by models such as Sumformer (Alberti et al., 2023), FluRKA (Gupta et al., 2023), Scatterbrain (Chen et al., 2021a), Low-Rank Transformer (LRT) (Winata et al., 2020), Performer (Choromanski et al., 2021), Random Feature Attention (RFA) (Peng et al., 2021), Linear Transformer (Katharopoulos et al., 2020), and Linformer (Wang et al., 2020) enhance the efficiency by utilizing low-rank representations of the self-attention matrix or by adopting attention kernelization techniques. Specifically, low-rank methods focus on compacting the dimensions of attention keys and values. For example, Linformer (Wang et al., 2020) proposes to segment scaled dot-product attention into smaller units via linear projection. Kernelization, a variant of low-rank techniques, focuses on approximating the attention matrix (Choromanski et al., 2020). For example, Performer (Choromanski et al., 2021) condenses softmax attention-kernels using positive orthogonal random features, outperforming Reformer and Linformer on long protein sequence benchmark. Sumformer (Alberti et al., 2023) approximates the equivariant sequence-to-sequence function, offering a universal solution for Linformer and Performer.
@@
Fixed Pattern Strategies. Fixed pattern strategies adopted by models such as (Pagliardini et al., 2023), Big Bird (Zaheer et al., 2020), Poolingformer (Zhang et al., 2021), Longformer (Beltagy et al., 2020), Blockwise Transformer (Qiu et al., 2020), and Sparse Transformer (Child et al., 2019) improve efficiency by sparsifying the attention matrix. This is achieved by confining the attention scope to predetermined patterns, such as local windows or fixed-stride block patterns. For instance, the attention mechanism adopted by Longformer (Beltagy et al., 2020), designed as an alternative to conventional self-attention, merges local windowed attention with globally oriented attention tailored to specific tasks. Pagliardini et al. (2023) expand FlashAttention (Dao et al., 2022) to support a broad spectrum of attention sparsity patterns, including key-query dropping and hashing-based attention techniques, achieving a multi-fold runtime speedup on top of FlashAttention on long text benchmark.
@@
Learnable Pattern Strategies. Different from fixed pattern strategies, learnable pattern strategies adopted by models such as HyperAttention (Han et al., 2024), Reformer (Kitaev et al., 2020), Sparse Sinkhorn Attention (Tay et al., 2020), Clustered Attention (Vyas et al., 2020), ClusterFormer (Wang et al., 2022b), and Routing Transformer (Roy et al., 2021) improve efficiency by learning token relevance and subsequently grouping tokens into buckets or clusters. As an example, HyperAttention (Han et al., 2024) proposes a parameterization for spectral approximation and employs two key metrics: the maximal column norm in the normalized attention matrix and the row norm ratio in the unnormalized matrix after large entry removal. It also utilizes the learnable sort locality-sensitive hashing (sortLSH) technique and fast matrix multiplication via row norm sampling. The experiment results show that HyperAttention enhances both inference and training speeds for LLMs with only minimal performance degradation, giving significant speed improvements compared to FlashAttention (Dao et al., 2022) on long contexts.
Hardware-Assisted Attention. Hardware-assisted attention focuses on developing hardware-specific techniques to enhance attention efficiency. For example, FlashAttention (Dao et al., 2022) reduces the number of memory access between GPU high-bandwidth memory (HBM) and GPU on-chip SRAM when calculating the attention module in LLMs. Instead of transmitting the values and results between HBM and SRAM multiple times as is done in the standard attention mechanism, FlashAttention combines all the attention operations into one kernel and tiles the weight matrices into smaller blocks to better fit the small SRAM as shown in Figure 14. As a result, only one communication is required to process each attention block, which significantly enhances the efficiency for processing the entire attention block. vAttention (Prabhu et al., 2024) is proposed to store KV cache in contiguous virtual memory without committing physical memory ahead-of-time. It avoids the software complexity to store KV cache by leveraging CUDA support of low-level virtual memory APIs.
2.5.2 Mixture of Experts (MoE)